My rectangular cabbage patch has a total area of 115 sq. ft. Find its perimeter in terms of the width x.
Answer by Alan3354(69443)

My rectangular cabbage patch has a total area of 115 sq. ft. Find its perimeter in terms of the width x.
--------------
P = 2L + 2x --- L = Length
x*L = 115 --> L = 115/x
---
P = 2x + 230/x